# Neural Context Flows for Meta-Learning of Dynamical Systems

### v4

@@ 5.4 Conclusion and Future Work
Acknowledgments and Disclosure of Funding
This work was supported by UK Research and Innovation grant EP/S022937/1: Interactive Artificial Intelligence, EPSRC program grant EP/R006768/1: Digital twins for improved dynamic design, EPSRC grant EP/X039137/1: The GW4 Isambard Tier-2 service for advanced computer architectures, and EPSRC grant EP/T022205/1: JADE: Joint Academic Data science Endeavour-2.
@@ Appendix A Algorithms & Proofs
A.1 Proof of Proposition 3.1
For ease of demonstration, we propose an equivalent formulation of Proposition 3.1 that disregards the first input of $f$ not necessary for its proof. Although we will consistently write terms like $f(x)$, we emphasize that $x$ is meant to stand for the context, not the state variable.
Proposition A.1 (Second-order Taylor expansion with JVPs).
Assume $f:\mathbb{R}^{d_{\xi}}\rightarrow\mathbb{R}^{d}$ is $\mathcal{C}^{2}$. Let $x\in\mathbb{R}^{\xi}$, and define $g:y\mapsto\nabla f(y)(x-y)$. The second-order Taylor expansion of $f$ around any $x_{0}\in\mathbb{R}^{d_{\xi}}$ is then expressed as
@@
Plugging this into Equation 12, we have
@@
While this expression of the second-order Taylor expansion of a vector-valued function makes its implementation memory-efficient via Automatic Differentiation (AD), it still relies on nested derivatives, which scale exponentially with the order of the Taylor expansion due to avoidable recomputations. For even higher-order Taylor expansions, this scaling is frightfully inefficient. Taylor-Mode AD is a promising avenue to address this issue Bettencourt et al. [2019].
@@ A.2 Identifiability of physical parameters
Our work critically builds on CAMEL Blanke and Lelarge [2024] which extensively studies a model similar to NCF-$\mathit{t}_{1}$ like a linearly parametrized system. Its Proposition 1 states, informally, that in the limit of vanishing training loss $\mathcal{L}(\cdot,\cdot,\cdot)=0$, the relationship between the learned context vectors and the system parameters is linear and can be estimated using ordinary least square. This forces the model to learn a meaningful representation of the system instead of overfitting the examples from the training tasks. We aim to validate this proposition in our setting by adapting the LV experiment.
@@
For CoDA Kirchmeyer et al. [2022] and NCF-$\mathit{t}_{1}$, we use the exact same network architecture: a 4-layer MLP with 224 hidden units. This means no context nor state network is used in NCF-$\mathit{t}_{1}$: the context vector of size $d_{\xi}=2$ is directly concatenated to the state vector as done by Zintgraf et al. [2019]. We note that the results obtained using this configuration further indicate the superiority of NCF, even when model comparison centers on their main/root networks (see also Park et al. [2023] for a similar model comparison based on parameter count) [^5] As reported in Appendix B, the primary model comparison approach in this work counts all learnable parameters, including hypernetworks if involved. Only the context vectors are exempt from this count..
@@
Upon meta-training and meta-testing, we set out to recover the underlying parameters of the Lotka-Volterra systems via a linear transformation of the learned context vectors. We fit a linear regression model to the 9 meta-training context vectors, using the true physical parameters as supervision signal. We test on the 4 adaptation contexts. The results, displayed in Figure 6, adequately illustrate interpretability as stated in [Blanke and Lelarge, 2024, Proposition 1]. They show that our trained meta-parameters recover the underlying system parameters up to a linear transform, and thus enable zero-shot (physical parameter-induced) adaptation via inverse regression.
@@ A.3 Convergence of proximal alternating minimization
For clarity of exposition, Theorem 3.2 expressing the convergence of Algorithm 1 to second-order stationary points is repeated below.
Theorem A.2 (Convergence to second-order stationary points).
@@
We use this section to briefly emphasize that the assumptions of KL property Attouch et al. [2010] and Lipschitz continuity are mild and easily achievable with neural networks. Interestingly, boundedness of the sequence $(\theta_{k},\xi^{1:m}_{k})$ is guaranteed a priory if $\mathcal{L}$ is coercive Li et al. [2019], a property we encouraged by regularizing $\ell$ wrt weights and contexts as in Equation 9.
@@ A.4 Ordinary alternating minimization
Here we provide an additional procedure (Algorithm 3) for training Neural Context Flows which we reserved for the NCF-$\mathit{t}_{1}$ variant. While stronger assumptions Li et al. [2019] are needed to establish convergence guarantees like its proximal extension, is it relatively easier to implement, and exposes fewer hyperparameters to tune.
@@ A.5 Bulk adaptation algorithm
    Slower convergence: the bulk algorithm could take longer to converge to poorer context vectors when the jointly adapted environments are all very far apart. This is because the contextual self-modulation process would be rendered useless by such task discrepancy, and sampling $j=e$ in the context pool $\mathrm{P}$ to return to standard a Neural ODE in Section 3 would be harder. We find in practice that manually setting $j=e$ by adjusting the vector field to forego the Taylor expansion works even better than the sequential counterpart (see Figure 3). A more direct way of achieving the same result is to disregard $\mathrm{P}$ and retain $f_{\theta}$ (rather than its Taylor expansion) in Equation 5.
@@
A.6 Scalability of the NCF algorithms
@@
Primarily, NCF’s training process can be distributed and parallelized across environments and trajectories (see Equation 10 and Algorithm 4), a feature that distinguishes it from baseline methods which are limited to parallelization across trajectories. Furthermore, the framework employs an efficient approach, as outlined in Proposition 3.1, to avoid materializing Jacobians or Hessians wrt potentially large context vectors, thereby significantly enhancing scalability. Additionally, NCF utilizes only first-order gradients wrt model weights $\theta$, in contrast to methods like CAVIA that require second-order information in their bi-level optimization loop.
@@
Scalability can also be evaluated in terms of component size, particularly in meta-learning adaptation rules that incorporate additional contextual parameters beyond shared model weights. These components may include encoders, hypernetworks, and other mechanisms for generating or refining context vectors necessary for task adaptation. The complexity and memory requirements of contextual meta-learning rules, which are directly related to the size of these components, can be quantified by their parameter count among other metrics. In this regard, NCF maintains a constant memory cost $O(1)$, while baseline methods such as CAVIA and CoDA require additional memory to produce better contexts (see [Park et al., 2023, Table 1]).
@@
Despite these advantages, the NCF framework may face challenges related to memory and computational efficiency due to the requirement of solving $p$ Neural ODEs in (3), as opposed to a single one. In scenarios where all training environments are utilized (i.e., $p=m$), this results in a quadratic cost $O(m^{2})$ for Algorithms 1 and 3. However, our ablation studies in Section D.1 demonstrate that competitive performance can be achieved on most problems using as few as $p=2$ neighboring environments. Additional studies in Sections D.2 and D.3 establish the necessity of expressive context vectors Voynov and Babenko [2020] and validate the efficacy of the 3-networks architecture, respectively.

### v5

@@ 5.3 Conclusion and Future Work
Ethics Statement
While the benefits of NCFs are evidenced in Section 5.1, its negative impacts should not be neglected. For instance, malicious deployment of such adaptable models in scenarios they were not designed for could lead to serious adverse outcomes. With that in mind, our code, data, and models are openly available at https://github.com/ddrous/ncflow.
Acknowledgments
This work was supported by UK Research and Innovation grant EP/S022937/1: Interactive Artificial Intelligence, EPSRC program grant EP/R006768/1: Digital twins for improved dynamic design, EPSRC grant EP/X039137/1: The GW4 Isambard Tier-2 service for advanced computer architectures, and EPSRC grant EP/T022205/1: JADE: Joint Academic Data science Endeavour-2. We extend our sincere gratitude to the anonymous reviewers whose insightful feedback and rigorous commentary substantially enhanced the quality and clarity of this manuscript. We also acknowledge Amarpal Sahota for valuable discussions on uncertainty quantification methodologies.
Supplementary Material for Neural Context Flows for Meta-Learning of Dynamical Systems
@@ Appendix A Algorithms & Proofs
A.1 Second-order Taylor expansion with JVPs
For ease of demonstration, we propose an equivalent formulation of Proposition 1 that disregards the first input of $f$ not necessary for its proof. Although we will consistently write terms like $f(x)$, we emphasize that $x$ is meant to stand for the context, not the state variable.
Proposition 1 (Second-order Taylor expansion with JVPs).
Assume $f:\mathbb{R}^{d_{\xi}}\rightarrow\mathbb{R}^{d}$ is $\mathcal{C}^{2}$. Let $x\in\mathbb{R}^{d_{\xi}}$, and define $g:y\mapsto\nabla f(y)(x-y)$. The second-order Taylor expansion of $f$ around any $x_{0}\in\mathbb{R}^{d_{\xi}}$ is then expressed as
@@ A.1 Second-order Taylor expansion with JVPs
Plugging this into Eq. 12, we have
@@
While this expression of the second-order Taylor expansion of a vector-valued function makes its implementation memory-efficient via Automatic Differentiation (AD), it still relies on nested derivatives, which scale exponentially with the order of the Taylor expansion due to avoidable recomputations. For even higher-order Taylor expansions, this scaling is frightfully inefficient. Taylor-Mode AD is a promising avenue to address this issue Bettencourt et al. (2019).
@@ A.2 Identifiability of physical parameters
Identifying the underlying physical parameters of a system is of enduring interest to scientific machine learning practitioners. Our work critically builds on CAMEL Blanke & Lelarge (2024) which extensively studies a model similar to NCF-$\mathit{t}_{1}$ as a linearly parametrized system. Its Proposition 1 states, informally, that in the limit of vanishing training loss $\mathcal{L}(\cdot,\cdot,\cdot)=0$, the relationship between the learned context vectors and the system parameters is linear and can be estimated using ordinary least square. This forces the model to learn a meaningful representation of the system instead of overfitting the examples from the training tasks.
We now reformulate Blanke & Lelarge (2024)’s identifiability result for linear systems trained with first-order Taylor expansion and the Random-All pool-filling strategy (see Section A.6), then we provide an alternative proof suited to our setting. Building on notations from Eqs. 1 and 4, we restate that our goal is to approximate the true vector field based on
$\frac{\text{d}x}{\text{d}t}(t)=f_{\text{true}}(x(t),c),\qquad\text{and }\qquad\frac{\text{d}x}{\text{d}t}(t)=f_{\theta}(x(t),\xi),\qquad\forall t\in\left[0,T\right].$
We drop the dependence on $x\in\mathbb{R}^{d}$ to ease notations in the sequel (the $\nabla$ will henceforth indicate gradients wrt $\xi$). As such, the predictor $f_{\theta}:\mathbb{R}^{d_{\xi}}\rightarrow\mathbb{R}^{d}$ is parametrized as a neural network and learned, while $f_{\text{true}}:\mathbb{R}^{d_{c}}\rightarrow\mathbb{R}^{d}$ is known and affine, i.e. $\exists P\in\mathbb{R}^{d\times d_{c}}\text{ and }p\in\mathbb{R}^{d}$ such that $f_{\text{true}}(c)=Pc+p$.
Proposition 2 (Identifiability of affine systems).
Assume $d_{\xi}\geq d_{c}$, that $P$ is full-rank, and that $f_{\textup{true}}$ is differentiable. In the limit of zero training loss in Eq. 10, $f_{\theta}$ is affine on an open region of $\mathbb{R}^{d_{\xi}}$. Furthermore, there exists $Q\in\mathbb{R}^{d_{c}\times d_{\xi}}$ and $q\in\mathbb{R}^{d_{c}}$ such that for any meta-trained $\xi\in\{\xi^{e}\}_{e=1}^{m}$ and its corresponding underlying parameter $c\in\{c^{e}\}_{e=1}^{m}$, we have $c=Q\xi+q$.
Proof.
Let $e\in\operatorname{[\![}1,m\operatorname{]\!]}$. In the limit of zero training loss, $f_{\theta}$ coincides with its first-order Taylor expansion in a neighborhood $U(\xi^{e})$ which contains all other $\{\xi^{j}\}_{j=1}^{m}$. We can write
$f_{\theta}(\xi)=f_{\theta}(\xi^{e})+A(\xi-\xi^{e}),\qquad\forall\,\xi\in U(\xi^{e})$ (14)
where $A=\nabla f_{\theta}(\xi^{e})$ is constant.
Similarly, for $j\in\operatorname{[\![}1,m\operatorname{]\!]}$, there exists an open set $U(\xi^{j})$ which contains all other $\{\xi^{e}\}_{e=1}^{m}$, such that
$f_{\theta}(\xi)=f_{\theta}(\xi^{j})+B(\xi-\xi^{j}),\qquad\forall\,\xi\in U(\xi^{j})$ (15)
where $B=\nabla f_{\theta}(\xi^{j})$ is constant.
To show that necessarily $A=B$, let’s consider without loss of generality $\xi\in U(\xi^{e})\cap U(\xi^{j})$ (which is non-empty since both $\xi^{e}$ and $\xi^{j}$ are included in both sets) and proceed by contradiction, assuming $A\neq B$. Let $v\in\mathbb{R}^{d_{\xi}}$ sufficiently small, we can set $\xi=\xi^{e}+tv$ in Eq. 14 and write the directional derivative
$\lim_{t\to 0}\frac{f_{\theta}(\xi^{e}+tv)-f_{\theta}(\xi^{e})}{t}$ $=\lim_{t\to 0}\frac{f_{\theta}(\xi^{e})+A(\xi^{e}+tv-\xi^{e})-f_{\theta}(\xi^{e})}{t}$
$=\lim_{t\to 0}\frac{tAv}{t}$
$=Av.$ (16)
We also write, using Eq. 15, the same directional derivative as
$\lim_{t\to 0}\frac{f_{\theta}(\xi^{e}+tv)-f_{\theta}(\xi^{e})}{t}$ $=\lim_{t\to 0}\frac{f_{\theta}(\xi^{j})+B(\xi^{e}+tv-\xi^{j})-f_{\theta}(\xi^{e})}{t}$
[…]
$=Bv.$ (17)
For $v\notin\ker{(B-A)}$, we have $Av\neq Bv$ which contradicts with the uniqueness of directional derivatives for differentiable functions Spivak (1965). This shows that $f_{\theta}$ is affine on the open set $U=\bigcap\limits_{e=1}^{m}U(\xi^{e})$.
Furthermore, using Eq. 14, we have
$f_{\theta}(\xi)-A\xi=f_{\theta}(\xi^{e})-A\xi^{e},\qquad\forall\,\xi\in U$ (18)
which is valid for all $e\in\operatorname{[\![}1,m\operatorname{]\!]}$. This indicates that the right hand side of Eq. 18 is constant, i.e. $\exists\,\tilde{q}\in\mathbb{R}^{d_{\xi}}$ such that $\forall\,e\in\operatorname{[\![}1,m\operatorname{]\!]}$, $f_{\theta}(\xi^{e})-A\xi^{e}=\tilde{q}$ We then use the fact that in the limit of zero training loss, the predicted and true vector fields coincide for a context $\xi\in U$ and its corresponding underlying parameters $c$ :
$f_{\text{true}}(c)=f_{\theta}(\xi)\Rightarrow Pc+p$ $=A(\xi-\xi^{e})+f_{\theta}(\xi^{e})\qquad\text{for}\,e\in\operatorname{[\![}1,m\operatorname{]\!]}$
$=A\xi+\tilde{q}.$
Since $d_{\xi}\geq d_{c}$ and $P$ is full-rank, its rows are linearly independent, guaranteeing the existence of a pseudo-inverse. We can thus write $c=Q\xi+q$ with
$Q=(P^{T}P)^{-1}P^{T}A,\qquad\text{ and }\qquad q=(P^{T}P)^{-1}P^{T}[\tilde{q}-p].$ (19)
The closed form expression Eq. 19 can be challenging to derive, especially since $P$ and $p$ might not be fully known when collecting data. So similar to Blanke & Lelarge (2024), one can perform post-training, ordinary least squares regression on observed $\{c^{e}\}_{e=1}^{m^{\prime}}$ (with $m^{\prime}\leq m$) to estimate the optimal $Q^{*}$ and $q^{*}$ :
$Q^{*},q^{*}\in\underset{Q,q}{\text{argmin}}\frac{1}{2}\sum_{e=1}^{m^{\prime}}\|Q\xi^{e}+q-c^{e}\|^{2}_{2}.$ (20)
Experimental validation.
@@
We validate Proposition 2 by modifying the Lotka-Volterra (LV) experiment. For CoDA Kirchmeyer et al. (2022) and NCF-$\mathit{t}_{1}$, we use the exact same network architecture: a 4-layer MLP with 224 hidden units. This means no context nor state network is used in NCF-$\mathit{t}_{1}$: the context vector of size $d_{\xi}=2$ is directly concatenated to the state vector as done by Zintgraf et al. (2019). We note that the results obtained using this configuration further indicate the superiority of NCF, even when model comparison centers on their main/root networks (see also Park et al. (2023) for a similar model comparison based on parameter count) [^5] As reported in Appendix D, the primary model comparison approach in this work counts all learnable parameters, including hypernetworks if involved. Only the context vectors are exempt from this count..
@@
After meta-training and meta-testing, we set out to recover the underlying parameters of the Lotka-Volterra systems via a linear transformation of the learned context vectors. We fit a linear regression model to the 9 meta-training context vectors, using the true physical parameters as supervision signal. We test on the 4 adaptation contexts. The results, displayed in Fig. 6, adequately illustrate interpretability as stated in (Blanke & Lelarge, 2024, Proposition 1) and our Proposition 2. They show that our trained meta-parameters recover the underlying system parameters up to a linear transform, and thus enable zero-shot (physical parameter-induced) adaptation via inverse regression.
Robustness to noise.
Additionally, system identification with NCFs is robust to noise in the trajectory. We show this empirically by corrupting the single trajectory in each adaptation environment with a Gaussian noise scaled by a factor of $\eta$. Upon addition of this noise, sequential adaptation is performed to recover new $\xi$ which are then transformed into $c$ and plotted. The weight $Q$ and bias $q$ of the affine transform are fitted on the training environments and their corresponding underlying parameters, which are unchanged across all noise levels. Fig. 8 shows that the reconstruction [^6] Sample context vectors pre-reconstruction can be observed in Fig. 21 as the structure is preserved with changing seeds. MSE remains low despite the noise (compared to CoDA’s MSE of $1.57\times 10^{-2}$ when $\eta=0$ shown in Fig. 6), and the physical system remains visually identifiable, especially in the convex hull of training environments. Outside the convex hull, the identifiability is notably worse with $\eta\geq 0.1$ indicating that this noise level is excessively high given the range of the LV state values.
Figure 8: Robustness of NCF-$\mathit{t}_{1}$ when identifying physical parameters with varying levels of noise $\eta$ injected into the adaptation trajectory. The MSE is reported with the standard deviation across 10 runs with different seeds for the Gaussian noise. (Left) $\eta=0.01$, (Middle) $\eta=0.05$, (Right) $\eta=0.1$.
@@ A.3 Convergence of proximal alternating minimization
For clarity of exposition, Theorem 1 expressing the convergence of Algorithm 1 to second-order stationary points is repeated below.
Theorem 1 (Convergence to second-order stationary points).
@@
We use this section to briefly emphasize that the assumptions of KL property Attouch et al. (2010) and Lipschitz continuity are mild and easily achievable with neural networks. Interestingly, boundedness of the sequence $(\theta_{k},\xi^{1:m}_{k})$ is guaranteed a priory if $\mathcal{L}$ is coercive Li et al. (2019), a property we encouraged by regularizing $\ell$ wrt weights and contexts as in Eq. 9.
@@ A.4 Ordinary alternating minimization
Here we provide an additional procedure (Algorithm 3) for training Neural Context Flows which we reserved for the NCF-$\mathit{t}_{1}$ variant. While stronger assumptions Li et al. (2019) are needed to establish convergence guarantees like its proximal extension, is it relatively easier to implement, and exposes fewer hyperparameters to tune.
@@ A.5 Bulk adaptation algorithm
    Slower convergence: the bulk algorithm could take longer to converge to poorer context vectors when the jointly adapted environments are all very far apart. This is because the contextual self-modulation process would be rendered useless by such task discrepancy, and sampling $j=e$ in the context pool $\mathrm{P}$ to return to standard a Neural ODE in Section 3 would be harder. We find in practice that manually setting $j=e$ by adjusting the vector field to forego the Taylor expansion works well (see Fig. 3). A more direct way of achieving the same result is to disregard $\mathrm{P}$ and retain $f_{\theta}$ (rather than its Taylor expansion) in Eq. 5.
@@
A.6 What’s in a context pool ?
The content of the context pool $\mathrm{P}$ not only defines the candidate trajectories we get from Eq. 8, but also the speed and memory cost of the meta-training process. Based on intuitive understanding of the role of $\mathrm{P}$, we outline 3 tunable pool-filling strategies for selecting $p$ neighboring contexts:
  • Random-All (RA): all $p$ distinct contexts can be selected by randomly drawing their indices from $\operatorname{[\![}1,m\operatorname{]\!]}$. By repeatedly doing so, we maximize long-range interactions to provide the broadest form of self-modulation – since information can always (in the stochastic limit) flow from any environment into $e$.
  • Nearest-First (NF): only the $p$ closest contexts to $\xi^{e}$ are selected, thus encouraging environments to form clusters. (Note, however, that if $p=1$, then $\mathrm{P}=\{e\}$ itself, and no self-modulation occurs.) In one ablation study, we observe that this strategy is the most balanced with regard to training time and performance (see Section C.3).
  • Smallest-First (SF): the smallest contexts in $L^{1}$ norm are selected first. Since an environment with context close to 0 can be interpreted as an environment-agnostic feature like in Kirchmeyer et al. (2022); Blanke & Lelarge (2024), this strategy prioritizes the flow of information from that base or canonical environment to the one of interest $e$.
A.7 Scalability of the NCF algorithms
@@
Primarily, NCF’s training process can be distributed and parallelized across environments and trajectories (see Eq. 10, Fig. 13, and Algorithm 4), a feature that distinguishes it from baseline methods which are limited to parallelization across trajectories. Furthermore, the framework employs an efficient approach, as outlined in Proposition 3.1, to avoid materializing Jacobians or Hessians wrt potentially large context vectors, thereby significantly enhancing scalability. Additionally, NCF utilizes only first-order gradients wrt model weights $\theta$, in contrast to methods like CAVIA that require second-order information in their bi-level optimization loop.
@@
Scalability can also be evaluated in terms of component size, particularly in meta-learning adaptation rules that incorporate additional contextual parameters beyond shared model weights. These components may include encoders, hypernetworks, and other mechanisms for generating or refining context vectors necessary for task adaptation. The complexity and memory requirements of contextual meta-learning rules, which are directly related to the size of these components, can be quantified by their parameter count among other metrics. In this regard, NCF maintains a constant memory cost $O(1)$, while baseline methods such as CAVIA and CoDA require additional memory to produce better contexts (see (Park et al., 2023, Table 1)).
@@
Despite these advantages, the NCF framework may face challenges related to memory and computational efficiency due to the requirement of solving $p$ Neural ODEs in equation 3, as opposed to a single one. In scenarios where all training environments are utilized (i.e., $p=m$), this results in a quadratic cost $O(m^{2})$ for Algorithms 1 and 3. However, our ablation studies in Section C.1 demonstrate that competitive performance can be achieved on most problems using as few as $p=2$ neighboring environments. Additional studies in Sections C.2 and C.4 establish the necessity of expressive context vectors Voynov & Babenko (2020) and validate the efficacy of the 3-networks architecture, respectively.